I need to learn how to put the following equation
the perimeter of a rectangle is 30 feet. the length is 3 feet less than twice the width. what are the dimensions of the rectangle?

Answer by checkley77(12844)

P=2L+2W
L=2W-3
30=2(2W-3)+2W
30=4W-6+2W
30=6W-6
6W=30+6
6W=36
W=36/6
W=6 ANS.
L=2*6-3
L=12-3=9 ANS.
pROOF:
30=2*9+2*6
30=18+12
30=30